Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 296     Entry time: 10 Aug 2023, 07:22
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, up to 181 pulses at 564 kHz, ~1.8 mJ. Overall stable delivery.

Optical laser delivery

  • 800 nm, 50 fs, 564 kHz, up to 185 pulses. Overall stable delivery.

Achievements/Observations

  • Tested first implementation of new automatic jet alignment device.
  • Setup of photon arrival time monitor.

Issues

  • No analog output from SPB_MKB_MVE/ACTRL/AX_FINE. Solved by EEE via DOC.
  • Issue with MKBs coordinated motions under investigation by CTRL via DOC. Nothing obviously wrong, investigation still ongoing.
  • SPB_OHD_SCR/MOTOR/ZOOM and SPB_OHD_SCR/MOTOR/FOCUS not working. Some parameters lost during summer maintenance period. Fixed by DOC/EEE.
  • SPB_IRU_INJMOV/MOTOR/Y regulargly going into error state.
  • No access to key XFEL/DESY services (elog, Redmine, etc.) for a while.
  • SPB optics hutch shutter opened during the night, around 2:10, when we were not present causing a beam delivery interruption of about 5 minutes in SASE1/3.
ELOG V3.1.4-7c3fd00